OpenStudy (anonymous):

How many formula units make up 11.2 g of magnesium chloride MgCl2? how do i set this up?

OpenStudy (anonymous):

find molar mass of MgCL2 = 95.211g 11.2g/95.211g = .1176 * 6.02 * 10^23 formula units = 1.95 * 10^21 molecules